OpenStudy (anonymous):

find the fiexed points g(x)=x^2+8x+6

OpenStudy (anonymous):

fixed points sorry

OpenStudy (zarkon):

solve \[g(x)=x\]

OpenStudy (anonymous):

\[g \left( x \right)=x ^{2}+8x+6\] \[x ^{2}+8x+6=0\]

OpenStudy (anonymous):

Is this right so far?

OpenStudy (zarkon):

\[x^2+8x+6=x\] solve for x

OpenStudy (anonymous):

so what do I do from here

OpenStudy (anonymous):

I always get confused when I have to solve for x

OpenStudy (anonymous):

\[\left( x+4 \right)\left( x+2 \right)=?\]

OpenStudy (zarkon):

are you just guessing?

OpenStudy (anonymous):

no I must be wrong

OpenStudy (zarkon):

\[x^2+8x+6=x\] \[x^2+7x+6=0\]

OpenStudy (anonymous):

why did the 8 change to a 7?

OpenStudy (zarkon):

I subtracted x from both sides

OpenStudy (anonymous):

ok that would make sence

OpenStudy (zarkon):

now factor

OpenStudy (anonymous):

do I go \[\left( x+6 \right)\left( x+1 \right)\]

OpenStudy (anonymous):

oh wait like this\[\left( x+1 \right)\left( x+6 \right)=x ^{2}+7x+6x\]

OpenStudy (zarkon):

so what is x equal to?

OpenStudy (anonymous):

x=1 x=6

OpenStudy (zarkon):

no

OpenStudy (anonymous):

oh boy

OpenStudy (anonymous):

so shat did I do wrong

OpenStudy (zarkon):

you factored correctly...but those are not the values of x

OpenStudy (zarkon):

if x+1=0 then x=-1

OpenStudy (anonymous):

how do I know that?

OpenStudy (zarkon):

subtract 1 from both sides

OpenStudy (zarkon):

\[\left( x+6 \right)\left( x+1 \right)=0\]

OpenStudy (anonymous):

ok so I have a fixed point of x=-1

OpenStudy (zarkon):

so what is the other one?

OpenStudy (anonymous):

I will be honest I am not sure

OpenStudy (anonymous):

will it be x+6=0 then x=-6

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

yes yea I think I jst might be getting this

OpenStudy (anonymous):

can you walk me through one with a square root I will post

OpenStudy (zarkon):

I will be away from my computer for about an hour or so...I'll help you later if no one else does.

OpenStudy (anonymous):

ok thanks so much You help me a ton